Dr. Dawson is considering two business opportunities. Both require an initial investment of $200,000. The first will return $50,000 at the end of each of the next 6 years, while the second will return $35,000 at the end of each of the next 10 years. Calculate the present value of the profit from these two businesses at each of the following discount rates: 7%, 8%, 9%, 10%, and 12%. Use the graph that follows to plot the relationship between the present value (PV) of profit from each of the two businesses and the interest rate. At approximately what interest rate would Dr. Dawson be indifferent between the two businesses? Over what range of interest would be Buisness 1 be preferred? Over what range of interst rates would be Business 2 be prferred?
1
Expert's answer
2019-09-20T09:27:38-0400

NPV1(7%) = -200,000 + 50,000/1.07 + 50,000/1.07^2 + 50,000/1.07^3 + 50,000/1.07^4 + 50,000/1.07^5 + 50,000/1.07^6 = 38,326.98.

NPV1(8%) = 31,143.98.

NPV1(9%) = 24,295.93.

NPV1(10%) = 17,763.03.

NPV1(12%) = 5,570.37.

NPV1(7%) = -200,000 + 35,000/1.07^1 + ... + 35,000/1.07^10 = 45,825.35.

NPV1(8%) = 34,852.85.

NPV1(9%) = 24,618.02.

NPV1(10%) = 15,059.85.

NPV1(12%) = -2,242.19.

Dr. Dawson would be indifferent between the two businesses at approximately 9%.

Buisness 1 would be between 9% and 12%.

Buisness 2 would be between 7% and 9%.